# E2-BKI: Evidential Ellipsoidal Bayesian Kernel Inference for Uncertainty-aware Gaussian Semantic Mapping

### v1

@@ VIII-C Gaussian as an Ellipsoid
We choose $\tau$ to enclose a specified fraction $p$ of the Gaussian’s probability mass. For a $d$-dimensional Gaussian, the squared Mahalanobis distance follows a chi-squared distribution with $d$ degrees of freedom, so $\tau$ is the chi-squared quantile at level $p$:
$\tau=F^{-1}_{\chi^{2}_{d}}(p),$ (29)
where $d=3$ for 3D voxel mapping and $d=2$ for 2D BEV mapping. Intuitively, $p$ sets how much of the distribution the ellipsoid contains (e.g., $p=0.10$ encloses $10\%$).
The distance $\mathbf{d}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j})$ is then defined as the shortest projection from the query $\mathbf{\hat{x}}_{m}$ onto the ellipsoidal surface:
$\min_{\mathbf{v}\in\mathbb{R}^{3}}\|\mathbf{\hat{x}}_{m}-\mathbf{v}\|^{2}\text{ s.t. }(\mathbf{v}-\bm{\mu}_{j})^{\top}\bm{\Sigma}_{j}^{-1}(\mathbf{v}-\bm{\mu}_{j})=\tau.$ (30)
We compute this projection by reducing it to a single scalar parameter. After translating and rotating the query into the ellipsoid’s principal-axes frame, the closest point is obtained by contracting the rotated coordinates toward the center along each axis; a single scalar governs all contractions and is chosen so that the point lies exactly on the ellipsoid boundary. Rotating and translating back then gives the closest point $\mathbf{v}^{\star}$, and we define $\mathbf{d}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j}):=\|\mathbf{\hat{x}}_{m}-\mathbf{v}^{\star}\|$. If the query $\mathbf{\hat{x}}_{m}$ lies inside the ellipsoid $\mathcal{E}_{j}$, we set $\mathbf{d}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j})=0$.
@@ VIII-D Uncertainty Decomposition
While our formulation in (9) enhances continuous semantic mapping, reliable uncertainty estimation $u_{m}$ for the posterior $\bm{\alpha}_{m}$ remains crucial. A common surrogate is the normalized Dirichlet variance $\mathrm{Var}[\hat{\bm{\theta}}_{m}^{\psi_{m}}]$, which monotonically decreases as evidence accumulates and can therefore underestimate uncertainty in semantically ambiguous regions. To mitigate overconfidence and provide reliable uncertainty estimates for safe decision-making, we decompose $u_{m}$ into semantic uncertainty $u_{m,\text{sem}}$ and evidence sparsity $u_{m,\text{spa}}$. Concretely, the semantic uncertainty is computed as a kernel-weighted average of per-primitive uncertainties $u_{j}$:
$u_{m,\text{sem}}=\frac{\sum_{j=1}^{J^{*}}\tilde{k}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j})\,u_{j}}{\sum_{j=1}^{J^{*}}\tilde{k}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j})}.$ (31)
@@
with $S_{m}=\sum_{c=1}^{C}\alpha^{c}_{0}+\sum_{j=1}^{J^{*}}\tilde{k}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j})$ from our update rule. Finally, the overall uncertainty is
@@ IX-A Datasets
OffRoad consists of RGB images ($2464{\times}1713$), Ouster OS1-128 LiDAR scans, and estimated robot poses [32] with manually annotated 2D semantic segmentation labels. LiDAR-camera projection utilizes an advanced extrinsic calibration procedure based on [33], ensuring accurate 2D-3D correspondence for reliable semantic mapping evaluation. Collected through high-speed navigation under diverse conditions including different seasons (spring, fall, and winter), lighting variations, and terrain types, this dataset presents significant challenges for reliable 3D mapping. The dataset comprises 8,8938{,}893 images for 2D semantic segmentation network training and 1,4001{,}400 frames ($300$, $400$, $300$, $400$ per sequence) for 3D semantic mapping evaluation.
@@ IX-B Experimental Setup
3D Semantic Mapping We adopt a single set of mapping hyperparameters across all datasets as summarized in Table I. The only dataset-specific setting is the K-Means++ cluster count for initializing Gaussian primitives: 256 for RELLIS-3D and 1,024 for the other datasets, to balance map compactness and geometric detail.
@@
Evaluation Since dense ground truth occupancy maps are rarely available in real-world datasets and are expensive to create, we build the map using every fifth frame and evaluate against reference point cloud queries constructed from the full sequence, which reflects online operation under limited sensing. Accordingly, the $20\%$ data availability in Fig. 6 represents our primary evaluation setting. To construct the reference queries, we accumulate all frames, voxelize the aggregated cloud at one quarter of the evaluation resolution ($0.05$ m), and retain only voxels whose points exhibit consistent semantics.
@@
For evaluation, semantic quality is reported with classwise Intersection over Union ($\mathrm{IoU}$) and its mean ($\mathrm{mIoU}$), and we additionally report an accuracy metric ($\mathrm{Acc}$) that couples geometric occupancy with semantic correctness. For uncertainty calibration we use the Brier Score ($\mathrm{BS}$ $\downarrow$). Let $\text{TP}_{c},\text{FP}_{c}$, and $\text{FN}_{c}$, represent the number of true positive, false positive, and false negative queries, respectively. Then, metrics are computed as:
$\text{$\mathrm{IoU}$}_{c}$ $=\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FP}_{c}+\text{FN}_{c}},\quad\text{$\mathrm{mIoU}$}=\frac{1}{C}\sum_{c=1}^{C}\text{$\mathrm{IoU}$}_{c},$
$\mathrm{Acc}$ $=\frac{1}{Q}\sum_{c=1}^{C}\text{TP}_{c},\quad\mathrm{BS}=\frac{1}{Q^{\prime}}\sum^{Q^{\prime}}_{m=1}(\mathbb{1}_{\text{correct}}-\mathrm{Conf}_{m})^{2},$ (34)
@@
where $Q$ is the total number of queries, $Q^{\prime}$ is the number of occupied queries, and $\mathbb{1}_{\text{correct}}$ is assigned $1$ if the semantic label is correctly estimated as the ground truth and 0 otherwise. The confidence for query $m$ is $\mathrm{Conf}_{m}=1-\widehat{\mathrm{Var}}[\hat{\theta}_{m}^{\psi}]$ for variance-based baselines (normalized from (13)), and $\mathrm{Conf}_{m}=1-u_{m}$ for ours in (33).
IX-C Supplementary Results
@@
BEV Semantic Mapping Fig. 11 illustrates qualitative results of BEV semantic mapping experiments. On RELLIS-3D, our method effectively suppresses noisy vegetation predictions and yields cleaner boundaries along roads and traversable ground. On OffRoad, it preserves thin structures and produces sharp guardrail predictions that baselines often miss. On KITTI-360, it reconstructs a more complete and consistent road surface even with sparse mapping input.
Figure 9: Uncertainty map comparison for the third-row scene of Fig. 5.
Ablation Studies We conduct ablation studies on RELLIS-3D as summarized in Table VII. Starting from the S-BKI baseline, EBS addresses semantic uncertainty and demonstrates substantial improvements. The introduction of isotropic Gaussian primitives transitions the framework to primitive-based operations, allowing local observations to be aggregated into semantically coherent units. We also evaluate Farthest Point Sampling (FPS), but K-Means++ yields more balanced clusters and further stabilizes performance. Extending primitives to anisotropic Gaussians enables geometry-aligned kernels, mitigating spatial uncertainty. Local merging improves mapping stability and computational efficiency without sacrificing accuracy, while adaptive pruning significantly enhances semantic performance. Finally, uncertainty decomposition in (33) improves posterior calibration, resulting in more interpretable and reliable uncertainty estimates, as shown in Fig. 9.
@@
TABLE VII: Ablation study evaluating the contribution of each component in our framework, alongside comparisons with ConvBKI variants.

### v2

@@ VIII-C Gaussian as an Ellipsoid (Section IV-D)
We choose $\tau$ to enclose a specified fraction $\tilde{\tau}$ (we use $10\%$) of the Gaussian’s probability mass. For a $d$-dimensional Gaussian, the squared Mahalanobis distance follows a chi-squared distribution with $d$ degrees of freedom, so $\tau$ is the chi-squared quantile at level $\tilde{\tau}$:
[…]
where $d=3$ for 3D voxel mapping and $d=2$ for 2D BEV mapping. Intuitively, $\tilde{\tau}$ specifies the fraction of the Gaussian probability mass enclosed by the ellipsoid (e.g., $\tilde{\tau}=0.10$ encloses $10\%$ of the mass).
The distance between a primitive $\mathcal{G}_{j}$ and a query point $\mathbf{\hat{x}}_{m}$, denoted by $d^{\ast}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j}):=d^{\ast}$, is defined as the shortest Euclidean distance from $\mathbf{\hat{x}}_{m}$ to the ellipsoidal surface of $\mathcal{E}_{j}$. The closest point $\mathbf{v}^{\ast}$ on the surface is obtained by solving the following constrained minimization:
$d^{\ast}=\min_{\mathbf{v}\in\mathbb{R}^{3}}\|\mathbf{\hat{x}}_{m}-\mathbf{v}\|^{2}\text{ s.t. }(\mathbf{v}-\bm{\mu}_{j})^{\top}\bm{\Sigma}_{j}^{-1}(\mathbf{v}-\bm{\mu}_{j})=\tau,$ (31)
with $d^{\ast}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j})=0$ when $\mathbf{\hat{x}}_{m}\in\mathcal{E}_{j}$.
@@ VIII-D Uncertainty Decomposition
While our formulation in (9) enhances continuous semantic mapping, reliable uncertainty estimation $u_{m}$ for the posterior $\bm{\alpha}_{m}$ remains crucial. A common surrogate is the normalized Dirichlet variance $\mathrm{Var}[\hat{\bm{\theta}}_{m}^{\psi_{m}}]$, which monotonically decreases as evidence accumulates and can therefore underestimate uncertainty in semantically ambiguous regions. To mitigate overconfidence and provide reliable uncertainty estimates for safe decision-making, we decompose $u_{m}$ into semantic uncertainty $u_{m,\text{sem}}$ and evidence sparsity $u_{m,\text{spa}}$.
Concretely, the semantic uncertainty is computed as a kernel-weighted average of per-primitive uncertainties $u_{j}$:
$u_{m,\text{sem}}=\frac{\sum_{j=1}^{J^{*}}\tilde{k}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j})\cdot u_{j}}{\sum_{j=1}^{J^{*}}\tilde{k}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j})}.$ (32)
@@
with $S_{m}=\sum_{c=1}^{C}\alpha^{c}_{0}+\sum_{j=1}^{J^{*}}\tilde{k}(\mathbf{\hat{x}}_{m},\mathcal{G}_{j})$ from our update rule. Finally, the overall uncertainty $u_{m}$ is obtained by combining the semantic and spatial uncertainty components:
@@ IX-A Datasets
OffRoad consists of RGB images ($2464{\times}1713$), Ouster OS1-128 LiDAR scans, and estimated robot poses [29] with manually annotated 2D semantic segmentation labels. LiDAR-camera projection utilizes an advanced extrinsic calibration procedure based on [19], ensuring accurate 2D-3D correspondence for reliable semantic mapping evaluation. Collected through high-speed navigation under diverse conditions, including different seasons (spring, fall, and winter), lighting variations, and terrain types, this dataset presents significant challenges for reliable 3D mapping. The dataset comprises 8,8938{,}893 images for 2D semantic segmentation network training and 1,4001{,}400 frames ($300$, $400$, $300$, $400$ per sequence) for 3D semantic mapping evaluation.
@@ IX-B Implementation Details
3D Semantic Mapping We adopt a single set of mapping hyperparameters across all datasets as summarized in Table I. The only dataset-specific setting is the number of K-Means++ clusters for initializing Gaussian primitives: 256 for RELLIS-3D and 1,024 for the other datasets, to balance map compactness and geometric detail.
IX-C Evaluation Setup
@@
Dense ground-truth occupancy maps are rarely available in real-world datasets and are costly to construct. To reflect online operation under limited sensing, we therefore build maps using every fifth frame and evaluate them against reference point-cloud queries constructed from the full sequence. Accordingly, the $20\%$ data availability in Fig. 6 represents our primary evaluation setting. To construct the ground-truth queries, we accumulate all frames, voxelize the aggregated cloud at one quarter of the evaluation resolution ($0.05$ m), and retain only voxels whose points exhibit consistent semantics.
@@
Evaluation Metrics For evaluation, semantic quality is reported with class-wise Intersection over Union ($\mathrm{IoU}$) and its mean ($\mathrm{mIoU}$), and we additionally report an accuracy metric ($\mathrm{Acc}$) that couples geometric occupancy with semantic correctness. Let $\text{TP}_{c},\text{FP}_{c}$, and $\text{FN}_{c}$, represent the number of true positive, false positive, and false negative queries for class $c$, respectively. Then, semantic quality metrics are computed as follows:
$\text{$\mathrm{IoU}$}_{c}$ $=\frac{\text{TP}_{c}}{\text{TP}_{c}+\text{FP}_{c}+\text{FN}_{c}},\quad\text{$\mathrm{mIoU}$}=\frac{1}{C}\sum_{c=1}^{C}\text{$\mathrm{IoU}$}_{c},$ (35)
$\mathrm{Acc}$ $=\frac{1}{Q}\sum_{c=1}^{C}\text{TP}_{c},$ (36)
where $Q$ is the total number of ground-truth queries. For uncertainty calibration we use the Brier Score ($\mathrm{BS}$ $\downarrow$) and the Expected Calibration Error ($\mathrm{ECE}$ $\downarrow$):
$\mathrm{BS}$ $=\frac{1}{Q^{\prime}}\sum^{Q^{\prime}}_{m=1}(\mathbb{1}_{\text{correct},m}-\mathrm{Conf}_{m})^{2},$
$\mathrm{ECE}$ $=\sum^{B}_{b=1}\frac{|\mathcal{B}_{b}|}{Q^{\prime}}|\mathrm{Acc}(\mathcal{B}_{b})-\mathrm{Conf}(\mathcal{B}_{b})|,$ (37)
@@
where $Q^{\prime}$ is the number of occupied queries, and $\mathbb{1}_{\text{correct},m}$ is an indicator function that equals $1$ if the predicted semantic label of the $m$-th query matches the ground truth and 0 otherwise. The confidence of the $m$-th query is defined as $\mathrm{Conf}_{m}=1-\widehat{\mathrm{Var}}[\hat{\theta}_{m}^{\psi_{m}}]$ for variance-based baselines, where $\widehat{\mathrm{Var}}[\hat{\theta}_{m}^{\psi_{m}}]$ denotes the normalized variance obtained from $\mathrm{Var}[\hat{\theta}_{m}^{\psi_{m}}]$ in (13). For our method, the confidence is given by $\mathrm{Conf}_{m}=1-u_{m}$, as defined in (34).
[…]
Figure 9: Uncertainty map comparison for the fourth-row scene of Fig. 5.
X Supplementary Results
X-A Quantitative Results
We present more complete quantitative results corresponding to those reported in Table II and Table III. Since our Gaussian primitives naturally support continuous evaluation, we include continuous evaluation results. Moreover, we report $\mathrm{ECE}$ to provide a more comprehensive assessment of uncertainty calibration. The quantitative results are summarized in Table VII.
TABLE VII: Summary of semantic accuracy and uncertainty calibration on RELLIS-3D, OffRoad, and KITTI-360. Higher is better for $\mathrm{mIoU}$ and $\mathrm{Acc}$, lower is better for $\mathrm{BS}$ and $\mathrm{ECE}$.
RELLIS-3D OffRoad KITTI-360
Method
$\mathrm{mIoU}$ [%]
$\mathrm{Acc}$ [%]
$\mathrm{BS}$ $\downarrow$ [%]
$\mathrm{ECE}$ $\downarrow$ [%]
$\mathrm{mIoU}$ [%]
$\mathrm{Acc}$ [%]
$\mathrm{BS}$ $\downarrow$ [%]
$\mathrm{ECE}$ $\downarrow$ [%]
$\mathrm{mIoU}$ [%]
$\mathrm{Acc}$ [%]
$\mathrm{BS}$ $\downarrow$ [%]
$\mathrm{ECE}$ $\downarrow$ [%]
S-CSM 43.7 67.7 15.0 13.9 69.6 58.5 9.2 7.9 47.3 59.5 14.7 12.8
S-BKI 43.8 78.9 16.3 12.4 69.2 78.5 14.4 12.6 47.1 74.9 17.4 10.8
SEE-CSOM 44.6 80.0 14.1 11.6 68.8 80.3 8.7 4.8 46.9 76.1 13.4 8.2
ConvBKI 44.3 80.7 15.2 14.7 71.1 84.3 8.4 7.3 48.3 68.0 13.5 11.4
EBS 45.1 80.9 13.2 9.9 71.9 82.6 7.0 3.0 49.0 77.8 12.5 5.1
Ours 47.3 83.5 13.0 9.1 71.5 87.7 7.4 2.2 49.4 80.0 12.1 6.3
Ours (Cont.) 47.4 83.7 12.9 9.2 71.7 87.9 7.3 2.3 49.6 80.2 12.0 6.2
X-B Semantic BEV Mapping
@@
Fig. 11 illustrates qualitative results of semantic BEV mapping experiments. On RELLIS-3D, our method effectively suppresses noisy vegetation predictions and yields cleaner boundaries along roads and traversable ground. On OffRoad, it preserves thin structures and produces sharp guardrail predictions that baselines often miss. On KITTI-360, it reconstructs a more complete and consistent road surface even with sparse mapping input.
X-C Ablation Studies
@@
TABLE VIII: Ablation study evaluating the contribution of each component in our framework, alongside comparisons with ConvBKI variants.
@@
We conduct ablation studies on RELLIS-3D as summarized in Table VIII. Starting from the S-BKI baseline, EBS addresses semantic uncertainty and demonstrates substantial improvements. The introduction of isotropic Gaussian primitives transitions the framework to primitive-based operations, allowing local observations to be aggregated into semantically coherent units. We also evaluate Farthest Point Sampling (FPS), but K-Means++ yields more balanced clusters and further stabilizes performance. Extending primitives to anisotropic Gaussians enables geometry-aligned kernels, mitigating spatial uncertainty. Local merging improves mapping stability and computational efficiency without sacrificing accuracy, while adaptive pruning significantly enhances semantic performance. Finally, uncertainty decomposition in (34) improves posterior calibration, resulting in more reliable uncertainty estimates, as shown in Fig. 9.
X-D Sensitivity Analysis
We conduct a sensitivity analysis to evaluate the robustness of E2-BKI with respect to its hyperparameters. E2-BKI involves the following hyperparameters: $\alpha_{0}^{c}$, $\ell$, $\beta$, $\tilde{u}$, $\epsilon$, $d_{L}$, $d_{S}$ (Table I), and $\tau$ (30). Following prior work [7, 17], we fix the Dirichlet prior $\alpha_{0}^{c}=0.001$ and set the kernel length-scale $\ell$ equal to the voxel resolution ($0.2\mathrm{m}$).
Uncertainty Threshold The dynamic uncertainty threshold $\tilde{u}$ controls whether an incoming semantic observation is integrated into the map based on its predicted uncertainty. This mechanism is adopted from [17] and is intended to suppress unreliable updates without introducing a fixed, dataset-specific cutoff. We analyze the effect of $\tilde{u}$ separately by explicitly comparing the proposed dynamic thresholding scheme against several static thresholds.
Concretely, we vary the uncertainty threshold over the range $[0.05,0.95]$ and evaluate both strategies. In the dynamic scheme, the hyperparameter $\tilde{u}$ specifies the proportion of the most uncertain primitives to be removed in each frame. For clarity, the horizontal axis in Fig. 12 is parameterized as $\texttt{Uncertainty Threshold}=1-\tilde{u}$. For example, a threshold value of $0.9$ corresponds to removing the top $10\%$ most uncertain primitives per frame in the dynamic scheme, whereas the static scheme removes all primitives whose uncertainty exceeds $0.9$ globally.
The resulting $\mathrm{Acc}$ and $\mathrm{mIoU}$ are summarized in Fig. 12. As the threshold decreases, fewer primitives are retained, leading to degraded $\mathrm{Acc}$ for both schemes. For $\mathrm{mIoU}$, the dynamic strategy achieves its best performance at a threshold of $0.9$ ($\tilde{u}=10\%$) and consistently outperforms any static threshold. This is because dynamic thresholding adapts to frame-wise variations in the uncertainty distribution, whereas a global static threshold is difficult to tune and may remove too few or too many primitives depending on the scene.
Based on these observations, we adopt the dynamic threshold with $\tilde{u}=10\%$ as our default. This setting provides stable performance and generalization across datasets without additional tuning, in contrast to EBS [17] which relies on dataset-specific values ($\tilde{u}=10\%$ on RELLIS-3D and $\tilde{u}=30\%$ on OffRoad). We note that when focusing on a single dataset or a specific downstream task, a carefully tuned static threshold may be beneficial. A hybrid strategy that combines dynamic thresholding with an additional static cutoff is a promising direction for future work.
Figure 12: Effect of the dynamic and static uncertainty threshold on $\mathrm{Acc}$ and $\mathrm{mIoU}$ on RELLIS-3D.
Additional Hyperparameters
Figure 13: Sensitivity analysis of the five hyperparameters used in E2-BKI on RELLIS-3D. For each hyperparameter, only that variable is varied while all others are fixed to the default configuration ($\beta=0.75,\ \epsilon=2.5,\ d_{L}=5.0\ell,\ d_{S}=\ell,\ \tilde{\tau}=0.1$). The definition of $\tilde{\tau}$ is provided in (30).
Figure 14: Sensitivity to the clustering strategy (K-Means++, FPS, and Random sampling) and the number of clusters.
Fig. 13 reports the effect of varying each hyperparameter while fixing others to the default values ($\beta=0.75,\epsilon=2.5,d_{L}=5.0\ell,d_{S}=\ell,\tilde{\tau}=0.1$) on RELLIS-3D. Both $\mathrm{Acc}$ and $\mathrm{mIoU}$ remain highly stable across a wide range of values. Notably, even with hyperparameter variations, our method consistently outperforms the best baseline results ($\mathrm{mIoU}$ =45.1=45.1 and $\mathrm{Acc}$ =80.9=80.9), indicating that the performance gain stems from the proposed Gaussian-primitive formulation rather than hyperparameter tuning. While larger $\beta$ improves accuracy, it increases inference latency. Thus, we select $\beta=0.75$ as a practical trade-off between accuracy and runtime.
A grid search over 2,0002{,}000 configurations revealed that $287$ settings ($14\%$) surpassed our default performance, with the best configuration reaching $\mathrm{Acc}$ =84.3=84.3 and $\mathrm{mIoU}$ =48.3=48.3. However, we deliberately employ the fixed default configuration throughout our experiments. This demonstrates that E2-BKI is robust to hyperparameter choices and generalizes well without requiring exhaustive, dataset-specific tuning.
X-E Clustering Strategy Analysis
While E2-BKI is theoretically agnostic to the clustering strategy used to form local point groups, the initialization quality can impact geometric fidelity and numerical stability. To examine the effect of different initialization schemes, we evaluate three approaches: K-Means++, Farthest Point Sampling (FPS), and Random sampling. Fig. 14 summarizes the quantitative results ($\mathrm{Acc}$, $\mathrm{mIoU}$, and $\mathrm{BS}$) on RELLIS-3D, OffRoad, and KITTI-360. While performance trends are consistent across strategies, K-Means++ demonstrates superior stability. FPS and Random sampling occasionally fail at high cluster counts (marked as ‘Failed’ in Fig. 14) due to insufficient points for valid covariance estimation, whereas K-Means++ remains stable.
As visualized in Fig. 15, the qualitative comparison on RELLIS-3D with only $64$ clusters reveals that K-Means++ preserves fine geometric details, whereas FPS and Random sampling exhibit artifacts such as holes or over-inflation. Based on these observations, we select K-Means++ as our default initialization strategy due to its stability. Regarding the cluster count, we employ $256$ clusters for RELLIS-3D, while scaling up to 1,0241{,}024 clusters for OffRoad and KITTI-360 to sufficiently capture the richer geometric details present in these environments.
Figure 15: Qualitative comparison of clustering strategies on RELLIS-3D using only $64$ clusters. The top row shows Gaussian primitives’ renderings, and the bottom row shows voxelized semantic maps. Red arrows highlight representative artifacts for FPS and Random sampling.
X-F Extended Runtime Analysis
We present an extended runtime analysis on RELLIS-3D in Table IX. All methods were evaluated on a laptop equipped with an Intel i7-12700H CPU and an NVIDIA RTX 3080 Ti Laptop GPU. The mapping speeds are given as the mean and standard deviation over five runs. While ConvBKI [34] attains the fastest runtime via GPU acceleration, it necessitates a substantial training overhead ($1759.5\pm 128.3$ seconds on RELLIS-3D). However, E2-BKI is training-free and could be further accelerated by adopting GPU acceleration.
TABLE IX: Mapping speed (Hz) comparison.
S-CSM S-BKI SEE-CSOM ConvBKI EBS Ours
3.44 $\pm$ 0.01 3.15 $\pm$ 0.04 3.00 $\pm$ 0.01 8.40 $\pm$ 0.21 1.65 $\pm$ 0.01 5.39 $\pm$ 0.11
Runtime Scalability
Figure 16: (a) Per-frame and (b) cumulative runtime over long-term sequences.
[…]
To evaluate the computational efficiency of E2-BKI in large-scale scenarios, we analyze its scalability relative to S-BKI [7] using a synthetic sequence of straight-line motion spanning 25,00025{,}000 frames. This setup simulates a significantly longer operation compared to the RELLIS-3D sequences used in the main experiments (avg. 1,0201{,}020 frames), thereby assessing the long-term behavior.
The results are presented in Fig. 16. As the trajectory continuously explores new environments, the number of Gaussian primitives $J$ grows linearly with the number of frames $F$. Since the current implementation manages primitives in a global set, the update cost scales with $J$, resulting in the observed quadratic growth ($\mathcal{O}(F^{2})$) in cumulative runtime (Fig. 16). Despite this, E2-BKI consistently maintains a lower computational cost than S-BKI. It is important to note that this straight-line motion represents a worst-case scenario for map expansion; in realistic deployments with revisited regions, merging and pruning mechanisms would effectively limit primitive growth. Furthermore, this quadratic scaling is an implementation artifact rather than a fundamental limitation. By introducing block-wise partitioning to bound the number of active primitives during updates, the system can achieve linear scalability $\mathcal{O}(F)$.
Impact of Voxel Resolution and Input Density
Figure 17: Runtime scalability of E2-BKI. The left (a, c) compare the total runtime of E2-BKI and S-BKI across voxel resolutions and relative input densities. The right (b, d) provides a component-wise breakdown of E2-BKI, separating Gaussian construction from voxelization.
[…]
We evaluate the computational scalability of E2-BKI by varying the voxel resolution from $0.4\mathrm{m}$ down to $0.01\mathrm{m}$ and adjusting the input point density. As illustrated in Fig. 17, S-BKI becomes prohibitively slow at fine resolutions ($0.05\mathrm{m}$), whereas E2-BKI remains tractable even at $0.01\mathrm{m}$. Notably, our method at $0.02\mathrm{m}$ operates faster than S-BKI at $0.1\mathrm{m}$, demonstrating superior efficiency. Furthermore, Fig. 17 confirms that E2-BKI scales linearly with input density. While S-BKI exhibits a steep increase in runtime, our framework grows at a moderate rate, reflecting natural data volume growth without significant algorithmic overhead.
The component-wise breakdown (Fig. 17) reveals that the Gaussian construction time is invariant to voxel resolution, confirming that our primitive generation is decoupled from the output grid. The exponential increase in runtime at extreme resolutions ($0.01\mathrm{m}$) is driven solely by the voxelization process, stemming from the inherent memory overhead of dense grids rather than the proposed Gaussian formulation. Regarding input density, Fig. 17 shows that Gaussian construction grows proportionally to the number of points, while voxelization remains relatively stable. This indicates that E2-BKI efficiently handles denser sensor streams.